OpenStudy (matimaticas):

logbase3(x-5) + logbase3(x+3) = 2

OpenStudy (misty1212):

HI!!

OpenStudy (matimaticas):

hello :)

OpenStudy (misty1212):

\[\log_3(x-5)+\log_3(x+3)=\log_3((x-5)(x+3))\] is a start

OpenStudy (matimaticas):

yeah I got stuck at that part

OpenStudy (misty1212):

oh then write in equivalent exponential form as \[(x-5)(x+3)=3^2\] and solve the resulting quadratic equation

OpenStudy (matimaticas):

do I FOIL